OpenStudy (anonymous):

what is the discriminant of x^2-12x+4=0 and what are the number and roots for the equation?

OpenStudy (anonymous):

my answer choices are: a:Discriminant is 144; the discriminant is greater than 0 and is the perfect square, the two roots are real and rational b: The discriminant is -128 because the discriminant is less than 0, the two roots are complex. c:The discriminant is 160. Beacuse the discriminant is greater than 0 and isn't a perfect square. d:The discriminant is -160. Because the discriminant is less than 0, the two roots are complex.

OpenStudy (anonymous):

\[\text{First, find the discriminant.}\\ \ \ \ \Delta = b^2-4ac\\ \ \ \ \ \ \ =(-12)^2-4(1)(4)\\ \ \ \ \ \ \ =144-16\\ \ \ \ \ \ \ =128\\ \text{The number of roots is the degree of the polynomial, i.e. 2.}\\ \text{Since }\Delta>0\text{ the polynomial has 2 distinct real roots.} \]

OpenStudy (anonymous):

thanks for helping me! I already found the answer to my problem. :)